# bar height=.6

### v1

@@
Computing is at a moment of profound opportunity and promise. Emerging applications – such as capable artificial intelligence, immersive virtual realities, and pervasive sensor systems – drive unprecedented demand for computer systems that offer scalable performance and environmental sustainability. Despite recent advances to achieve net zero carbon emissions, the computing industry’s gross energy usage continues to rise at an alarming rate, outpacing the growth of new energy installations and renewable energy deployments. Information and communication technology (ICT), all by itself, accounts for a shockingly large and rapidly growing slice of global greenhouse gas (GHG) emissions—estimates range from 2.1% to 3.9% (FREITAG2021100340, ).
@@
The International Telecommunication Union targets a 45% reduction in ICT emissions by 2030 (itu_carbon, ), in line with the Paris Agreement’s goal to limit warming to 1.5°C above pre-industrial levels. Satisfying demands for computing while meeting these goals will be difficult and expensive, demanding rigorous methods and solutions that balance sustainability benefits against implementation costs. To succeed, computer scientists, electrical engineers, environmental scientists, and economists must develop an ecosystem for sustainable computing with rigorous, transformative solutions to computing’s carbon problem, responding to the powerful call for action from Knowles et al. (10.1145/3503916, ): computing must end the “digital exceptionalism” that brushes aside its own carbon footprint because of the productivity and efficiency it provides to society.
@@
Embodied carbon from semiconductor manufacturing is a major contributor to emissions (KLINE2019322, ; gupta21, ; PIRSON2021128966, ), especially for mobile and embedded devices, due to high replacement rates and relatively low utilization. Nearly 75% of Apple’s corporate emissions are due to manufacturing (gupta21, ). Billions of devices are expected to come online by 2027, and their embodied carbon may approach one gigaton of CO2 per year, exceeding commercial aviation’s footprint (PIRSON2021128966, ). Moreover, the largest semiconductor fabrication companies consume electricity on par with hyperscale datacenters (Fig-2) and will consume even more for advanced technology nodes that require extreme ultraviolet (EUV) lithography. Embodied carbon costs increase even further when accounting for the gases required by semiconductor manufacturing (Fig-2).
@@
Operational carbon from powering computers is equally problematic. Over the next decade, annual ICT energy demand is projected to exceed 100 exajoules, reaching nearly 15% of the world’s energy production (src_decadal_plan, ). This explosive growth is driven by diverse applications such as AI, virtual spaces, Internet of Things, blockchains, etc. Electricity use at Google, Meta, and Microsoft grew at 25% per year from 2015 to 2021, nearly quadrupling over this period. In contrast, U.S. investments in renewable energy grew at only 7% per year (Fig-2). In 2021, hyperscale datacenters consumed an additional 19 TWh compared to 2020—nearly half of the 44 TWh of new renewable capacity that came online that year.
@@
Figure 4. Embodied carbon reduction target (45% by 2030) achieved with fab electricity growth/renewable scenarios and improving embodied carbon amortization via “reduce, reuse, recycle.”
@@
High percentage deployment of renewables presents its own set of challenges and opportunities. If renewable sources (primarily solar) account for more than 70% of electricity by the middle of the century, intra-day variability of energy generation will require changes to when and where computation might be performed. Sustainable datacenters must compute with carbon-free energy, delaying and boosting computation when supply is scarce and abundant, respectively, while still meeting strict demand response requirements. Otherwise, large renewable energy installations would have to be deactivated or curtailed during generation peaks with no matching consumption.
Our analysis of embodied carbon projects emissions from semiconductor fabrication in 2030. Fabs require large amounts of energy and incur high carbon costs. Although fabs could reduce their emissions by using carbon-free energy from renewable sources, at present, carbon-free electricity comprises a meager 6% of the total in Taiwan and South Korea, where a dominant fraction of chips are produced. Korea plans to raise the fraction of carbon-free energy to 20% of the total in 2030 and TSMC has made a corporate commitment to consume 40% renewable electricity in the same year (korea_renewables, ; tsmc_memory_website, ). Fig-4 presents three projections, considering what would happen if electricity consumption of fabs were to grow by 10% annually or 0%, and two scenarios for the fraction of renewables online by 2030, at 20% and 40%. Under status quo, even with 0% fab electricity growth, 20% renewables cannot reduce emissions by 45%, illustrated by the horizontal dotted line. This is partially due to gases (green portion) that comprise roughly 25% of total emissions but do not benefit from renewables (Fig-2). Reaching 45% reduction in carbon emissions by 2030 requires more aggressive and innovative measures.
[…]
A Call to Action. We envision coordinated research thrusts that produce design and management strategies for sustainable, next-generation computer systems. These strategies must seek to flatten and then reverse growth trajectories for computing power and carbon for society’s most rapidly growing applications such as artificial intelligence and virtual spaces (Section 1). We will require accurate models for carbon accounting and reporting in computing technology (Section 2). For embodied carbon, we must re-think conventional design strategies – over-provisioned monolithic servers, frequent hardware refresh cycles, custom silicon – and adopt life-cycle design strategies that more effectively reduce, reuse and recycle hardware at scale (Section 3). For operational carbon, we must not only embrace renewable energy but also manage systems to use that energy more efficiently (Section 4). Finally, new hardware design and management strategies must be cognizant of economic policy and regulatory landscape, aligning private initiatives with societal goals (Section 5). Many of these broader goals will require computer scientists to develop deep, enduring collaborations with researchers in economics, law, and industrial ecology to spark changes in broader practice.
1. Driving Applications and Studies
Growth in data volumes and model complexity have sparked debate about environmental sustainability. Perspectives on sustainable computing differ due to diverse assumptions, measurements, and models. Some argue that operational carbon dominates, calling for energy-efficient systems and demand response scheduling. Others argue that solutions for operational carbon are known and embodied carbon from manufacturing is the more pressing challenge. We view disparate perspectives as a call for better measurement and modeling. We also argue for a holistic set of solutions that balance operational and embodied carbon as any single solution exhibits diminishing returns.
Artificial Intelligence (AI). Advances in AI capability have been enabled by massively scaling deep models and their training data (ghorbani2022scaling, ; wei2022emergent, ), which in turn impact sustainability (strubell-etal-2019-energy, ; dodge22measuring, ). We need to benchmark AI’s carbon footprint for model development, training, and use. We must take an integrated hardware-software perspective as AI is in the midst of a hardware lottery: dominant models may not be fundamentally superior but rather are those that benefit most from hardware trends (hooker21, ). We must understand the net impact of custom hardware (e.g., TPU (jouppi17, )), which reduces operational carbon through energy efficiency but increases embodied carbon through semiconductor manufacturing.
AI must adapt to resource availability and system conditions. For sustainability, models must also adapt to diverse platforms, ensuring backward compatibility as we lengthen components’ refresh cycles and mitigate embodied carbon; we cannot simply deprecate models after three years and assume hardware updates. Models must adapt to platforms with varying degrees of programmability as architects transition from custom AI accelerators, for which fully amortizing embodied carbon is difficult, to more general ones. Finally, models must adapt for demand response, accounting for the availability of carbon-free compute across time.
[…]
Virtual Spaces. Future virtual spaces will integrate disparate technologies — artificial intelligence, gaming, social networks — and introduce computational demands that are greater than the sum of their already significant parts. AI is essential to existing social media and will become even more so. Recommendation systems will rely less on social networks and more on AI to discover and rank content (hazelwood18, ; wu19, ; gupta20, ). Generative AI will produce bots that serve as guides to the Internet. Enabling this experience at scale requires increasing throughput and decreasing latency for inference while managing carbon, a daunting task given the costs of today’s models.
Virtual spaces that provide immersive, richly rendered experiences will share attributes with cloud gaming, which evaluates game logic and renders frames using CPUs and GPUs in the datacenter. Cloud gaming requires high-resolution frames, high frame rates, and low latency. Virtual reality demands yet higher frame rates to avoid user disorientation. The gaming dimension produces interesting trade-offs between local and remote computation. Local compute on capable devices, such as VR headsets, increases embodied carbon, but remote compute at the datacenter increases operational carbon due to data movement. Shifting games from local compute on desktops, laptops, and consoles to remote compute in datacenters might increase total electricity use (summed at client and server) by up to 60%, 300%, and 200%, respectively (mills19, ).
2. Carbon Accounting
Effective carbon accounting is a prerequisite for sustainable computer system design and management but is exceptionally difficult for several reasons. First, embodied carbon models must account for increasingly complex semiconductor fabrication processes that are evolving for emerging technologies such as carbon nanotubes and photonic interconnect. Second, operational carbon must profile distributed energy use as computation draws on components that are physically distributed across machines. Most importantly, carbon accounting must be harmonized across the life cycle of computer systems and the industry’s many diverse participants, each with financial interests and intellectual property to guard.
Embodied Carbon. Computing’s embodied carbon is incurred during hardware manufacturing. We envision models that accommodate today’s increasingly heterogeneous systems as well as tomorrow’s circuits and devices. Models should account for emerging nanomaterials (hills18a, ; sabry_aly18, ; belmonte2020capacitor, ). They should include beyond-CMOS and photonic devices (pentecost2019maxnvm, ; wang2018integrated, ). And they should account for advanced integration (lau22_advanced_packaging, ; coudrain19_active_interposr, ; sabry_aly18, ).
The manufacture of “new” technologies actually leverage many existing fabrication process flows. By mixing and matching steps in mature flows—lithography, metal and oxide deposition, etching, thermal annealing, etc.—models can estimate embodied carbon for flows not yet in production. For example, the first monolithic 3D process flow that integrates two layers of next-generation transistors (1D semiconductors) and two layers of RRAM was recently deployed in a commercial foundry, SkyWater (srimani_and_hills20, ). The ”new” process flow consisted of existing steps (albeit in a different order) with the exception of one new step for depositing 1D semiconductors on the wafer’s upper layers.
Operational Carbon. Computing’s operational carbon depends on the amount and carbon intensity of the energy consumed. We must design energy profilers for individual tasks, helping operators track energy usage and guide management. Telemetry requires significant compute, network, and storage resources, especially when tracking tens of thousands of tasks at fine timescales. System telemetry will be combined with grid telemetry, which details renewable energy generation across time to estimate electricity’s carbon intensity. The marginal emission rate, which quantifies the carbon intensity of the most recently activated generation source on the grid’s dispatch stack, may overstate operational carbon because datacenters often negotiate purchase agreements and receive renewable energy credits from their investments in wind and solar generation.
Carbon Attribution. Telemetry lays foundations for attribution, which assigns responsibility for carbon to individual tasks in a datacenter. A task’s operational carbon depends on fine-grained energy telemetry and assessments of shared datacenter overheads (i.e., power usage effectiveness). A more subtle challenge arises when estimating a task’s share of embodied carbon for servers and datacenter infrastructure. Servers typically run co-located tasks and each task may use heterogeneous mixes of hardware that impact its share of embodied carbon. When a server idles, its embodied carbon is not associated with useful work yet must be attributed to tasks. Game theory and the Shapley value may provide frameworks for fair attribution (moulin04, ).
Accounting and Validation. We require reliable, harmonized, and transparent carbon accounting methods for computer systems. Energy use and its emissions are verifiable through the EPA’s carbon intensity statistics for power plants and directly measured energy for hardware components. Estimating energy use for fabrication is more difficult but can leverage published sustainability reports and datasets. Carbon from chemical/fuel combustion during fabrication benefits from a good understanding of combustion chemistry. Carbon accounting often leverages life cycle assessment (LCA) methodologies and data. Open-source carbon models would lay the foundations for engaging broad stakeholders and improving LCA models and results (sleep21, ). Such efforts are far behind in computing but other industries have harmonized accounting to reduce the uncertainty associated with LCA modeling choices. For example, the EPA and California both set standards to reduce life-cycle GHG emissions of fuels, using open-source LCA-based carbon accounting tools such as GREET (GREET, ) and OPGEE (ElHoujeiri2013OpensourceLT, ).
Risks. Accounting frameworks and models are only as good as the data they are applied to. Data for carbon intensity is supplied by public repositories for electricity grids (e.g., US EIA). This data could be combined with public traces of datacenter computation (e.g., Google (googletrace, ), Azure (microsofttrace, ), Alibaba (alibabatrace, )) to estimate operational carbon. Data for embodied carbon could be derived from first principles and from industrial datasets and publications. Open-sourcing carbon accounting frameworks creates transparency that may spur harmonized reporting from industry. Uncertainty will always exist in any model, but frameworks will show bounds for different possible futures.
3. Embodied Carbon Mitigation
Computing’s embodied carbon arises from the manufacture of systems. Chips have followed Moore’s Law, increasing chip functionality to lower costs of packaging and organizing chips into a system (moore65, ). Chips now integrate tens of billions of transistors for not only CPU cores but also domain-specific co-processors and controllers for memory, IO, and more. Although datacenters once used commodity servers (barroso03google, ), it now uses expensive servers, each with tens of CPU cores, terabytes of DRAM, and multiple GPUs with their own high-bandwidth memory. It has become impossible for a datacenter and its users to fully utilize these capabilities and amortize embodied carbon from hardware manufacturing.
Embodied carbon is often viewed through the lens of a waste-management hierarchy, which enumerates and ranks mitigation strategies (reducereuserecycle, ). For example, there are several strategies for purchasing a car, and these can be ranked from most to least sustainable: (1) avoiding the vehicle purchase entirely, (2) reusing or extending the life of an existing vehicle, (3) repairing a vehicle that includes recycled components from other disassembled vehicles. More generally, society-wide efforts seek a more circular economy (EllenMacArthurIntro, ; EllenMacArthurElectronics, ). Similar strategies are relevant to computing.
We must mitigate computation’s embodied carbon through interlocking reduce, reuse, and recycle strategies. Reducing hardware, decoupled systems will allow architects to manufacture, provision, and allocate precisely the mix of hardware required for application needs. Reusing hardware in these decoupled systems, operators can replace components based on individual technology advances or failure rates rather than based on the fastest evolving or least reliable component, thereby extending the hardware’s average tenure. Finally, Recycling hardware will require better instrumentation and health prediction to facilitate an efficient secondary market for components.
Reduce – Resource Disaggregation. We need modular design tools and strategies that decouple the provision of hardware functionality and tailor emerging technologies to reduce embodied carbon. Chip will benefit from heterogeneous integration, which manufactures chiplets with fine-grained functionality and combines them with fast interconnect (ucie_website, ; coudrain19_active_interposr, ). Chiplets are carbon-efficient because they manufacture precisely the required circuits and no more, reducing silicon area and improving manufacturing yield, which in turn reduces waste and carbon. Moreover, chiplets separate the manufacture of disparate capabilities—compute, memory, sensors—and use dedicated fabrication processes for each, reducing the number of process steps and associated carbon.
We also need datacenter-scale resource disaggregation, which organizes hardware into pools of network-attached components. Compute nodes offer many CPUs but little DRAM, whereas memory nodes offer the reverse. Disaggregation will require servers that can independently scale a specific hardware type. Heterogeneous integration will permit “lego-block” systems with custom core and memory configurations, better balancing the system and amortizing carbon. Such systems are more carbon-efficient than today’s servers, which provision many DRAMs for capacity but must also provision memory channels and processor sockets. For example, datacenter workloads often require high DRAM capacity but underutilize the bundled bandwidth and compute (malladi12, ).
Reuse – Lifespan Extension. Enabling component reuse improves sustainability by amortizing embodied carbon over a longer lifespan. Today, the typical server lifetime is three years, after which the entire rack is replaced with new hardware. Networking equipment lifetimes are longer, five years for switches/routers and ten years for the fiber cable plant, but periodic and wholesale replacement is still standard.
Recycle – Secondary Markets. A secondary market for hardware disassembles systems into constituent components and sells them for a second life, further amortizing embodied carbon. Transparency and data are needed for market efficiency. For instance, heavily used processors from hyperscale datacenters will have very different resale values than lightly used processors from enterprise deployments. Thus, data must be curated by manufacturers, sellers, or third parties so that consumers can intelligently assign value to pre-owned hardware. For statistical analysis, we will design frameworks for registering hardware components and reporting their usage. This would significantly expand the scope of economic activity for the semiconductor industry. For example, manufacturers might move toward leased equipment, which is known to be more highly utilized (and thus carbon-efficient) than owned equipment.
Risks. One might argue that reducing, reusing, and recycling hardware runs counter to manufacturers’ financial incentives. Manufacturers earn more revenue by selling more components, a major risk to sustainable design. Two factors mitigate this risk. Contrary to the historical commoditization of hardware components, modern datacenter operators are large customers that can demand and influence custom hardware and features, including those that enhance infrastructure efficiency and sustainability. Moreover, when successful, our research will produce a robust secondary market that expects reliability and provenance akin to the demand for cars with higher resale values and clear maintenance records.
4. Operational Carbon Mitigation
Operational carbon is determined by how computing consumes energy. Hyperscale datacenters consume tens of terawatt-hours of electricity, and associated operational carbon doubled from 2017 to 2020 due to continued datacenter construction (gupta21, ). Mitigating this carbon requires datacenters that compute with carbon-free energy, delaying and boosting computation when supply is scarce and abundant, respectively. But such schedules challenge conventional wisdom (fan07, ) in which datacenters compute constantly at peak power to amortize costs of building the facility and provisioning its power.
Measuring power and estimating operational carbon requires advances in system telemetry. How DR uses these modes should depend on concise signals of carbon intensity rather than the intricacies of energy markets. DR will require infrastructure such as servers, energy storage, and renewable energy generation; for example, Meta has invested in 5.8TW of renewable power and 180MWh of storage (vanlear21, ; meta21, ). We must also ensure reductions in operational carbon justify increases in embodied carbon.
@@ 4. Operational Carbon Mitigation
The datacenter will receive information about energy supply through concise interfaces. One interface would communicate real-time prices that incentivize datacenters to modulate energy use. This scenario would efficiently match supply and demand, but departs from today’s contracts that charge based on the amount of provisioned power rather than actual use. And what prices are required to achieve the desired demand response? An alternative interface communicates carbon intensity rather than price. But this scenario assumes datacenters would modulate demand to reduce operational carbon without compensation. For these and other interfaces, we will explore how datacenter power budgets vary across time to reduce operational carbon.
@@
Power Modulation. Each user must define and implement multiple operating modes that modulate power when required. Hardware mechanisms will rely on energy proportionality, the idea that power should rise and fall with workload. Energy-proportional hardware is difficult to design because most components have a significant fixed power cost (dissipated even at near-zero load). Decades of research have improved CPUs, but today’s datacenters deploy large memory systems and accelerators (GPUs, TPUs) that we will re-architect for energy proportionality. Memory will need new interfaces as today’s dissipate high fixed power to deliver high bandwidth. Accelerators will need better support for virtualization and resource sharing.
@@
Software mechanisms will rely on approximate, degraded computing. For online services, Internet applications implement contingency plans for site events, ensuring varying degrees of service that depend on system availability and downtime. We will explore real-time system design and anytime algorithms to provide a smoother spectrum of quality-power trade-offs than permitted in today’s systems. This approach generalizes the search engine’s strategy of delivering the most relevant results within some allotted time. For batch workloads, we will deepen our exploration of computational sprinting (fan16, ), which allow parallel workloads to dynamically consume additional resources as power budgets permit.

### v2

@@
Information and communication technology (ICT) accounts for a shockingly large share of global greenhouse gas (GHG) emissions—estimates range from 2.1% to 3.9% (FREITAG2021100340, ). To address this grand challenge, the International Telecommunication Union targets a 45% reduction in ICT emissions by 2030 (itu_carbon, ), in line with the Paris Agreement’s goal to limit warming to 1.5°C above pre-industrial levels. Satisfying demands for computing while meeting these goals will be difficult and expensive, demanding rigorous methods and solutions that balance sustainability benefits against implementation costs. To succeed, computer scientists, electrical engineers, environmental scientists, and economists must develop an ecosystem for sustainable computing with rigorous, transformative solutions to computing’s carbon problem, responding to the powerful call for action from Knowles et al. (10.1145/3503916, ): computing must end the “digital exceptionalism” that brushes aside its own carbon footprint because of the productivity and efficiency it provides to society.
We envision four broad research thrusts that are needed to produce design and management strategies for sustainable next-generation computer systems. First, we require accurate models for carbon accounting and reporting in computing technology. Second, for the embodied carbon emitted during the manufacture of hardware and infrastructure, we must adopt life-cycle design strategies that more effectively reduce, reuse and recycle hardware at scale. Third, for operational carbon associated with computing’s electricity use, we must not only embrace renewable energy but also manage systems to use that energy more efficiently. And fourth, we require integrated, cross-cutting strategies for hardware design and management because techniques that reduce operational carbon may increase embodied carbon and vice versa.
New hardware design and management strategies must be developed in context. These strategies must seek to flatten and then reverse growth trajectories for computing power and carbon for society’s most rapidly growing applications such as artificial intelligence. These strategies must also be cognizant of economic policy and regulatory landscape, aligning private initiatives with societal goals. Many of these broader goals will require computer scientists to develop deep, enduring collaborations with researchers in economics, law, and industrial ecology to spark changes in broader practice.
1. Driving Applications
Advances in artificial intelligence (AI) are enabled by massively scaling deep models and their training data (ghorbani2022scaling, ; wei2022emergent, ), which in turn impact sustainability (strubell-etal-2019-energy, ; dodge22measuring, ). Benchmarking AI’s carbon footprint for model development, training, and deployment could help researchers identify the most pressing challenges. An integrated hardware-software perspective will be particularly helpful as AI is in the midst of a hardware lottery: dominant models may be those that benefit most from hardware trends (hooker21, ). Researchers should explore the net impact of custom hardware, which could reduce operational carbon through energy efficiency but increase embodied carbon through semiconductor manufacturing.
The future of sustainable AI hinges on its ability to adapt in response to the varying availability of resources such as data, hardware, and electricity. First, there is a compelling need to design, train, and deploy AI models that offer performance, efficiency, and accuracy on a broad spectrum of hardware platforms. Such models would ensure backward compatibility for and equity of access to AI features, permitting users to slow the rate of hardware refreshes due to sustainability or financial constraints. Instead of deprecating systems after just a few years, how can we develop models and platforms that remain relevant over longer periods and better amortize the carbon costs of model training and deployment?
Second, there is a complementary need for programmable, reconfigurable hardware that support a broad spectrum of AI workloads. Such processors would allocate precisely the hardware required for data processing, training, or inference, consuming energy in proportion to utilization. Instead of designing static AI accelerators, how can we develop flexible, general processors that are relevant for large classes of AI computation and better amortize embodied carbon from semiconductor fabrication? Finally, demand response strategies would allow models and platforms to modulate their use of electricity based on its carbon intensity. To what extent can AI workloads be scheduled across time and space to reduce operational carbon?
If successful, this research agenda will reverse current trends and permit advanced AI with lower carbon costs. Google consumes 1.5-2.3TWh for AI, 10–15% of its total energy use (patterson22, ). Meta attributes 30% of its AI energy for data processing, 30% to model training, and 40% to inference (wu22, ). Studies for BLOOM’s 176B-parameter language model, a GPT-3 replica, are alarming. Training uses 433MWh and emits 25T-CO2e whereas inference uses 914KWh and emits 19kgs-CO2e per day assuming 558 requests per hour (luccioni2022estimating, ). Production models’ carbon footprints could easily be 1000x higher assuming one query from each of ChatGPT’s 13M daily unique visitors in January 2023 and considering interest in applications of this technology.
2. Carbon Accounting
Computing’s embodied carbon is incurred during hardware manufacturing. Modeling embodied carbon is exceptionally difficult because already complex semiconductor fabrication processes are evolving to accommodate emerging technologies such as nanomaterials (hills18a, ; sabry_aly18, ), photonic devices (wang2018integrated, ), and advanced heterogeneous integration (lau22_advanced_packaging, ; sabry_aly18, ). Yet we are optimistic because the manufacture of “new” technologies actually leverage many existing process flows. By mixing and matching steps in mature flows—lithography, metal and oxide deposition, etching, thermal annealing, etc.—we might estimate carbon for flows not yet in production. For example, the first monolithic 3D process flow that integrates next-generation transistors and RRAM was recently deployed in a commercial foundry, SkyWater (srimani_and_hills20, ). This ”new” flow re-orders existing steps and adds one new step for depositing 1D semiconductors.
Operational carbon depends on the amount and carbon intensity of the energy consumed. We must design energy profilers for individual tasks, helping operators track energy usage and guide management. System telemetry will be combined with grid telemetry, which details renewable energy generation across time. But estimating electricity’s carbon intesnity is non-trivial. The marginal emission rate, which depends on the most recently activated generation source on the grid, may overstate operational carbon because datacenters often negotiate purchase agreements and receive credits from their investments in renewable energy and because grid operators often transfer energy across boundaries of regional balancing authorities.
Telemetry lays foundations for attribution, which assigns responsibility for carbon to individual pieces of computation. A task’s operational carbon depends on fine-grained energy telemetry and allocation of shared datacenter overheads. Estimating a task’s share of embodied carbon for servers and datacenter infrastructure requires sophisticated analysis. Servers co-locate tasks and each task may use heterogeneous mixes of hardware that impact its share of embodied carbon. Game theory and the Shapley value may provide frameworks for fair attribution (llull17, ).
We require reliable, harmonized, and transparent carbon accounting methods. Energy use and its emissions are verifiable through the EPA’s carbon intensity statistics for power plants and directly measured energy for hardware components. Estimating energy use for fabrication is more difficult but can leverage published sustainability reports and datasets. Carbon from chemical and fuel combustion during fabrication benefits from a good understanding of the chemistry. Carbon accounting often leverages life cycle assessment (LCA) methods. Open-source models would lay the foundations for improving analysis and engaging stakeholders (sleep21, ). Such efforts are far behind in computing but other industries have harmonized accounting. For example, the EPA and California set standards to reduce GHG emissions from fuels, using open-source tools (GREET, ; ElHoujeiri2013OpensourceLT, ).
Risks. Accounting frameworks and models are only as good as the data they ingest. Data for embodied carbon could be derived from first principles and industrial datasets. Tools, such as ACT and imec.netzero, estimate yields from integrated circuit (IC) manufacturing, where industry data is closely guarded, using broad and parameterized models (e.g., Murphy, Poisson). Moreover, techniques in robust optimization can account for uncertainty from input data and models to produce feasible solutions with accompanying confidence intervals.
Validating embodied carbon models also present challenges. But we draw inspiration from the IC community, which has already developed a combined bottom-up, top-down approach for performance and power. Models at various levels – from simulating transistor physics to modeling datacenter power – shape our overall understanding of the industry. Similarly, we envision a standard approach to modeling embodied carbon. At the bottom, we could model energy used by individual fabrication process steps and machines. At the middle, we could model energy used and emissions produced by the fabrication facility. At the top, we could estimate overall production volumes and carbon footprints with published sustainability reports from industry leaders. As with IC models, we expect validation and accuracy to improve over time as more data become available and models are refined.
3. Embodied Carbon
@@
Embodied carbon from semiconductor manufacturing is a major contributor to emissions (KLINE2019322, ; gupta21, ; PIRSON2021128966, ), especially for mobile and embedded devices, due to high replacement rates and relatively low utilization. Nearly 75% of Apple’s corporate emissions are due to manufacturing (gupta21, ). Billions of devices are expected to come online by 2027, and their embodied carbon may approach one gigaton of CO2 per year, exceeding commercial aviation’s footprint (PIRSON2021128966, ). The largest semiconductor fabrication companies consume large amounts of electricity, especially for advanced technology nodes that require extreme ultraviolet lithography (Fig 2). Their carbon costs increase even further when accounting for the gasses required by semiconductor manufacturing.
Figure 2. Embodied carbon scenaris that vary fab electricity growth, renewable energy use, and 3R’s of circular economy.
Although fabs could reduce their emissions by using carbon-free energy from renewable sources, at present, carbon-free electricity is a meager 6% of the total in Taiwan and South Korea, where most chips are produced. TSMC and Korea plan to increase their use of carbon-free energy to 40% and 20% of their respective totals by 2030 (korea_renewables, ; tsmc_memory_website, ).
Fig-2 presents several scenarios for embodied carbon, varying growth in fab energy demand and renewable energy supply. Even under optimistic assumptions in which fab demand is unchanged (0%) and renewable energy supply increases by 20%, the industry will miss its goal of reducing emissions by 45%, which is illustrated by the dashed line. This outcome is partially explained by gases, which account for 25% of total emissions and are unaffected by the use of renewable energy. Thus, reducing embodied carbon by 45% requires more aggressive and innovative measures.
[…]
Reducing hardware, architects should explore system design strategies that manufacture, provision, and allocate precisely the mix of hardware required for application needs. We need modular design tools for heterogeneous integration, the idea that hardware functions can be designed and implemented separately as small chiplets and then connected with fast interconnect networks (coudrain19_active_interposr, ). Chiplets are more carbon-efficient as fabs would manufacture precisely the required circuits and no more, reducing silicon area and improving manufacturing yield, which in turn reduces waste and carbon. Moreover, fabs could separate the manufacture of disparate capabilities—compute, memory, sensors—and use dedicated process flows for each, reducing the number of process steps and associated carbon.
We also need datacenter-scale disaggregation, the idea that hardware could be organized into collections of network-attached components. Compute nodes would offer many CPUs but little DRAM, whereas memory nodes would offer the reverse. Disaggregation allows servers to independently scale a specific hardware type. ”Lego-block” systems with custom core and memory configurations would better balance the system and amortize carbon, but designing such systems and then managing them at scale, where heterogeneity creates complexity, remains difficult. Such systems are more carbon-efficient than today’s servers that inefficiently provision large quantitites of hardware in fixed proportions. For examples, today’s servers provision many DRAMs for capacity but must also inefficiently provision a corresponding number of memory channels and processor sockets even when workloads underutilize the bundled bandwidth and compute (malladi12, ).
Re-using hardware in these decoupled systems, operators might replace components based on individual technology advances or failure rates rather than based on the fastest evolving or least reliable component, thereby extending the hardware’s average tenure. Enabling component re-use improves sustainability by amortizing embodied carbon over a longer lifespan. Today, the typical server lifetime is three years, after which the entire rack is replaced with new hardware. Networking equipment lifetimes are longer, five years for switches/routers and ten years for the fiber cable plant, but periodic and wholesale replacement is still standard.
Recycling hardware will require better instrumentation and health prediction to facilitate an efficient secondary market that disassembles systems into constituent components and sells them for a second life, further amortizing embodied carbon. Transparency is needed for market efficiency. For instance, heavily used processors from hyperscale datacenters will have very different resale values than lightly used processors from enterprise deployments. Thus, data must be curated by manufacturers, sellers, or third parties so that consumers can intelligently assign value to pre-owned hardware. We will design frameworks for registering hardware components and reporting their usage for statistical analysis. This would significantly expand the scope of economic activity for the semiconductor industry. Manufacturers might move toward leased equipment, which is known to be more highly utilized (and thus carbon-efficient) than owned equipment.
We draw inspiration from the role that odometers, vehicle history reports, and certified pre-owned designations play in the secondary vehicle market. We envision hardware with “odometers” that account for their previous usage. The odometer will be implemented with dedicated, immutable, and tamper-resistant registers that count operations. A single odometer value is an imperfect proxy for history and additional measures are likely needed. For memories and disks, registers might count errors and faults in addition to read/writes. For all components, measures of physical conditions such as power variations, thermal stresses, and humidity will be helpful. Researchers will need to identify relevant features and develop compact representations, especially for longitudinal data such as temperature.
Risks. One might argue that reducing, re-using, and recycling hardware runs counter to manufacturers’ financial incentives. Manufacturers earn more revenue by selling more components, a major risk to sustainable design. Two factors mitigate this risk. Contrary to the historical commoditization of hardware components, modern datacenter operators are large customers that can demand and influence custom hardware and features, including those that enhance infrastructure efficiency and sustainability. Moreover, when successful, these research directions will produce a robust secondary market that expects reliability and provenance akin to the demand for cars with higher resale values and clear maintenance records.
4. Operational Carbon
@@
Over the next decade, annual ICT energy demand is projected to exceed 100 exajoules, reaching nearly 15% of the world’s energy production (src_decadal_plan, ). This explosive growth is driven by diverse applications such as artificial intelligence, virtual spaces, Internet of Things, blockchains, etc. Electricity use at Google, Meta, and Microsoft grew at 25% per year from 2015 to 2021, nearly quadrupling over this period. In contrast, U.S. investments in renewable energy grew at only 7% per year (Fig-4). In 2021, hyperscale datacenters consumed an additional 19 TWh compared to 2020—nearly half of the 44 TWh of new renewable capacity that came online that year.
@@
Demand Response. If renewable energy accounts for a large majority of the total by the middle of the century, intra-day supply variability will require changes to when and where computation is performed. Sustainable datacenters will need to delay and boost computation when carbon-free energy is scarce and abundant, respectively, while still meeting strict peformance requirements. Such schedules would require re-thinking conventional wisdom in which datacenters compute constantly at peak power to amortize costs of building the facility and provisioning its power (fan07, ).
Datacenters will need sophisticated demand response (DR) frameworks that modulate energy demand in response to carbon-free energy supply. DR will require coordination between the energy grid, datacenter operator, and datacenter users, and it will require hardware and software mechanisms that trade-off performance and power. Ideally, DR frameworks will both incentivize participation and guarantee service. Game theory might provide rigorous foundations for modeling and shaping system dynamics when users independently and selfishly pursue performance goals. Real-time scheduling and robust machine learning might ensure decisions satisfy diverse service obligations.
@@
The datacenter will receive information about energy supply through concise interfaces. One interface would communicate real-time prices that incentivize datacenters to modulate energy use. This scenario would efficiently match supply and demand, but departs from today’s contracts that charge based on the amount of provisioned power rather than actual use. And what prices are required to achieve the desired demand response? An alternative interface communicates carbon intensity rather than price. But this scenario assumes datacenters would modulate demand to reduce operational carbon without compensation.
@@
Power Modulation. Each user must define and implement multiple operating modes that modulate power when required. Hardware mechanisms will rely on energy proportionality, the idea that power should rise and fall with workload. Energy-proportional hardware is difficult to design because most components have a significant fixed power cost dissipated even at near-zero load. Decades of research have improved CPUs, but today’s datacenters deploy large memory systems and graphics processing units that will need to be designed for energy proportionality. Memory will need new interfaces as today’s dissipate high fixed power to deliver high bandwidth. Accelerators will need better support for virtualization and resource sharing, which better amortizes fixed power costs over more useful computation.
@@
Software mechanisms will rely on approximate, degraded computing. Online applications implement contingency plans for site events, ensuring varying degrees of service that depend on system availability and downtime. We will explore real-time system design and anytime algorithms to provide a smoother spectrum of trade-offs between quality and power than permitted in today’s systems. This approach generalizes the search engine’s strategy of delivering the most relevant results within some allotted time (reddi10, ). Strategies for computational sprinting allow workloads to dynamically consume additional resources as power budgets permit (fan16, ).